OpenStudy (anonymous):

solve using elimination: 4p+5q=7 3p-2p=34 im really confused with this stuff...

OpenStudy (anonymous):

i know how to do it, its just that i keep coming up with the wrong answers

OpenStudy (anonymous):

Is the second equation correct or did you mistake a p for a q?

OpenStudy (anonymous):

oh srry...the last equation is 3p-2q=34

hero (hero):

q = (7-4p)/5 q = (34-3p)/-2 q = q 5(34-3p) = -2(7-4p) 170 - 15p = -14 + 8p 184 = 23p p = 184/23 Good luck finding q

OpenStudy (anonymous):

i said elimintation

hero (hero):

That is elimination. I eliminated q from the equation and solved for p

OpenStudy (anonymous):

but u solved for q first.. i was taugtht to just multiply each side...

hero (hero):

I don't want to argue about it either

hero (hero):

Is it a crime to solve for q first? Nope

OpenStudy (anonymous):

srry didnt mean to offend u, im jus sooo confused i really hate these problems

OpenStudy (anonymous):

gimme a sex goodman.. I have to relearn how to do it real quick

OpenStudy (anonymous):

sec**

hero (hero):

Okay, I'll finish solving for q. Dragon fly might take forever

OpenStudy (anonymous):

okay, so an easier way is to jus solve for q?

OpenStudy (anonymous):

I can solve with subsitution.. I'm not great with elimination.

OpenStudy (anonymous):

same..erybody says elimination is the easiest but i dunt think so...i keep trying and end up with decimals when the answers are whole numbers....

hero (hero):

q = 115/23

hero (hero):

There's no easy way to solve for q

OpenStudy (anonymous):

the answer is 8,-5 according to the back of my textbook....

OpenStudy (anonymous):

I'm getting fractions/decimals..

OpenStudy (anonymous):

i swear i didnt..all ma answers come out as decimals too...and the answers come out to be hole nums too...

hero (hero):

I'll figure it out

OpenStudy (anonymous):

well I got q = 41/7

OpenStudy (anonymous):

thnxs now i dunt feel so stupid anymore for getiing the wrong answr cuz u too got wrong answrs...

OpenStudy (anonymous):

The back of the book could be wrong... Not sure it's been years since I did this.

OpenStudy (anonymous):

not all the problems could be wrong thou...

OpenStudy (anonymous):

Not all of them. Just that one. It's possible.. I'm trying

OpenStudy (anonymous):

well im getting all them as decimals or fractions...

OpenStudy (anonymous):

oh well ima try to do the rest of them...

OpenStudy (anonymous):

thnks for the help...

hero (hero):

Yeah, guess what? 184/23 = 8 and 115/23 = 5

OpenStudy (anonymous):

well it says her that its gotta be -5

hero (hero):

It is. I'm just showing you my calculations weren't that far off. You should be happy

OpenStudy (anonymous):

ezactly y i gave u all a medal...

hero (hero):

Actually I had -115/23, but when I put it into my calculator, I never use negative and I got excited, but as you can see I did get it. I just usually keep a fraction answer as a fraction.

hero (hero):

Dragonfly gives up too easily

OpenStudy (anonymous):

I haven't given up. Well I just did. But still I was trying to do it..

hero (hero):

I'm sure you saw the above. I had it the whole time, I just normally keep the answer in fraction form if a get a fraction. I had no idea that those fractions divided evenly into whole numbers

OpenStudy (anonymous):

i jus did it and i got the correct answer!!! im soo happy for ma self!!LOL

OpenStudy (anonymous):

after like a gaziillilon times thou...

hero (hero):

What method did you use?

hero (hero):

I only did it once. Everything I did after that was merely checking. I knew I had it

OpenStudy (anonymous):

elimintation..jus as i was asked to do...i jus multiplied the first equation by 2 and the second eq. by 5 and i got the answer!!!!

hero (hero):

Trust me....I did the exact same thing you did. You just don't understand my way yet. I bet your answer came out to a fraction like mine where you had to divide the fraction to get 8 and -5, correct?

OpenStudy (anonymous):

yes! I FEEL SOO SMART LOL!!

hero (hero):

Good for you

OpenStudy (anonymous):

now to do the other 50 problems like this...ughhhhh

hero (hero):

If they're exactly like that and you really do have fifty more, I feel sorry for you.

OpenStudy (anonymous):

yea wat ever...ma dumb teacher didnt even give us notes, jus said to read the example of them and stacks hmmk over us...

hero (hero):

Honestly, the reason I ended up doing these the way that I do it is because it is one way that works everytime. I don't have to try to decide whether I want to use substitution or elimination or addition/subtraction method. I have one method that I use the exact same way everytime and I get the right answer every time.

hero (hero):

trying to use all those different methods get too confusing

OpenStudy (anonymous):

Oh whatever. I was doing that too and it didn't turn out lol

OpenStudy (anonymous):

true..but i keep thinking wat if on an exam we were to do it a certain method and i didnt know how to do it? thats why i take the time to know them so i wont mess up.....

OpenStudy (anonymous):

I'm not even in school. I'm sitting here doing math for fun.

hero (hero):

If you have to solve by elimination or substitution, you could argue that my way is actually a combination of both.

hero (hero):

Dragonfly, I'm no different from you.

OpenStudy (anonymous):

It's been a while.. I just like doing it. I find it cool how I have a terrible memory but I can remember what I did in Algebra in hs.

OpenStudy (anonymous):

ughh u guys are soooo lucky, i got tons of hmmwk as a 9th grader in alg 2..IM SOO HAPPY I FOUND THIS SITE!!!

hero (hero):

One thing you'll notice about me is I don't use any of the traditional methods of Algebra to solve algebra problems...Never did.....(although I understand them) but I don't agree with them. I DON'T use FOIL, I don't use the QUAD FORM, I don't use SUB or Elimination methods

OpenStudy (anonymous):

It's alright. I help my sister with Algebra 2 almost daily. I use all of the traditional methods.. lol

OpenStudy (anonymous):

WOW.. im a dumb nut..i gotta do it old fashion style...:(

hero (hero):

None of those " traditional methods" are systematic. In fact, they're the reason why students become confused with Algebra

OpenStudy (anonymous):

woah..u should be a math prof. u might actually help a lot of ppl

hero (hero):

FOIL is more of an acronym than a truly algebraic method

OpenStudy (anonymous):

ughh i hate foil....

hero (hero):

Do you want to learn the other way to do it?

OpenStudy (anonymous):

how...

OpenStudy (anonymous):

FOIL means First Outer Inner Last

hero (hero):

Okay, suppose you had to multiply (2x+4)(3x+5) what would you do if you couldn't use FOIL?

OpenStudy (anonymous):

ya i noe i hate it i get lost in ma own math..

hero (hero):

dragonfly, I know what FOIL means

OpenStudy (anonymous):

hmm. our teacher told us that in that case we had to have everyitng memorized..

OpenStudy (anonymous):

That wasn't to you bright one.

hero (hero):

Dragonfly he knows what FOIL is too. He said he doesn't like it

OpenStudy (anonymous):

He said HOW. As in How is FOIL more of an acronym than algebraic method.

OpenStudy (anonymous):

I was going to elaborate on my posting but you both had moved on to another subject.

hero (hero):

Anyways, to multiply (2x+4)(3x+5) without FOIL , do this: